OpenStudy (anonymous):

write the equation 3x+2y=8 in slope intercept form

OpenStudy (anonymous):

so slope intercept form is y=mx+b so what you need to do is isolate the y from your equation. first subtract the 3x from both sides of the equation so you have -2y = -3x + 8 then divide both sides by -2 so your result is y = (3/2)x - 4 m is your slope, so slope is 3/2 and your intercept is b, so that is -4

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

slope intercept form is y=mx+b your question is 3x+2y=8 you want to isolate y, so start by subtracting 3x from both sides 2y=-3x+8 divide both sides by 2 y=-3/2x+8/2 reduce the fractions if you can y=(-3/2)x+4